I need help on this for a school project:
c=3.14d solved for d
a=3.14r^2 solve for r^2

Answer by checkley77(12844)

c=3.14d solved for d
d=c/3.14 ans.
a=3.14r^2 solve for r^2
r^2=a/3.14 ans.